### v1

@@ 3.2 Asymmetric case
$\mathcal{D}_{B,R}(\phi,\tau)<\mathcal{D}_{B,R}(\phi^{\prime},\tau)$ and $\mathcal{D}_{B,R}(\phi^{\prime},\tau^{\prime})<\mathcal{D}_{B,R}(\phi,\tau^{\prime}).$
Next, we add these inequalities:
$\mathcal{D}_{B,R}(\phi,\tau)+\mathcal{D}_{B,R}(\phi^{\prime},\tau^{\prime})<\mathcal{D}_{B,R}(\phi^{\prime},\tau)+\mathcal{D}_{B,R}(\phi,\tau^{\prime}).$
Next, we invoke the definition of $\mathcal{D}_{B,R}$ and have:
$\sum_{b\in B}\lvert b+\tau-\phi(b)\rvert+\lvert b+\tau^{\prime}-\phi^{\prime}(b)\rvert<\sum_{b\in B}\lvert b+\tau-\phi^{\prime}(b)\rvert+\lvert b+\tau^{\prime}-\phi(b)\rvert.$
This, in particular, means:
$0<\sum_{b\in B}\lvert b+\tau-\phi^{\prime}(b)\rvert+\lvert b+\tau^{\prime}-\phi(b)\rvert-\lvert b+\tau-\phi(b)\rvert-\lvert b+\tau^{\prime}-\phi^{\prime}(b)\rvert.$ (3.1)
We will show that this inequality is false for every term of this sum and derive a contradiction with the assumption that $\phi^{\prime}(b)<\phi(b)$. To that end, fix a term $b\in B$ and let $x\coloneqq b+\tau-\phi(b)$, $c_{1}\coloneqq\tau^{\prime}-\tau$, and $c_{2}\coloneqq\phi(b)-\phi^{\prime}(b)$. Observe, that for every $x\in\mathbb{R}$, when $c_{1},c_{2}>0$:
$0\geq\lvert x+c_{2}\rvert+\lvert x+c_{1}\rvert-\lvert x\rvert-\lvert x+c_{1}+c_{2}\rvert.$
This holds by case analysis on $x$. Therefore, each term of (3.1) has an opposite sign. Hence, (3.1) is false and we arrive at the contradiction. ∎
@@
A breakpoint $\tau$ of type (ii) corresponds to a pair of monotonically increasing matchings $\phi,\phi^{\prime}$ for which for all points $b\in B$ we have $\phi(b)\leq\phi^{\prime}(b)$. Furthermore, consider a run $b_{s},\ldots,b_{t}$ of $\phi$ and a point $b_{i}$ with $i\in\{s,\ldots,t\}$. If $\phi(b_{i})<\phi^{\prime}(b_{i})$, then $\phi(b_{j})<\phi^{\prime}(b_{j})$ for all $j\in\{i,\ldots,t\}$.
@@
In our sweep line algorithm we will maintain a current optimal matching $\phi$. At each breakpoint of type (i) we will have an event to update the cost function of the matching. Furthermore, it follows from Corollary 3.7 that when we sweep over a breakpoint of type (ii), we can decompose the changes to the matching using a series of atomic events. In each such atomic event there is some suffix $b_{j},\ldots,b_{t}$ of a run $b_{s},\ldots,b_{t}$ that $\phi$ currently matches to $r_{u-j},\dots,r_{u}$ that will become matched to $r_{u-j+1},\ldots,r_{u+1}$. As we argued in the proof of Lemma 3.8, the total number of such events is only $\mathcal{O}(nm)$. Next, we express how we can efficiently compute the next such atomic event, and handle it.
Consider a run $B_{i}=b_{s},\ldots,b_{t}$ induced by $\phi$ at time $\tau$. Our aim is to find the smallest $\tau^{\prime}\geq\tau$ at which there is an atomic type (ii) event involving a suffix $b_{j},\ldots,b_{t}$ of $B_{i}$. Hence, for a given suffix $b_{j},\ldots,b_{t}$, we wish to maintain when it starts being beneficial to match $b_{j},\ldots,b_{t}$ to $r_{u-j+1},\ldots,r_{u+1}$ rather than to $r_{u-j},\ldots,r_{u}$.
@@
Observe that this function is piecewise linear, and decreasing (more precisely, non-increasing). Moreover, the breakpoints coincide with type (i) breakpoints of $f$ at which $b+\tau$ coincides with a red point. Hence, in between any two consecutive events, we can consider $\Delta^{\prime}_{j}$ as a linear function. See Figure 3.4 for an illustration.
@@
Let $[\tau_{1},\tau^{\prime}]\ni\tau$ be a maximal interval on which $f(\tau)$ is linear, let $\tau^{\prime}$ be a type (ii) breakpoint, and let $\phi$ be an optimal matching for $\tau$. Then there is a run $B_{i}$ induced by $\phi$, and $\tau^{\prime}$ is a root of the function $g_{i}(\tau)$.
@@
At any moment of our sweep, we maintain a single piece of $g_{i}$. Hence, this piece is the lower envelope of a set of linear functions $\Delta_{s},\ldots,\Delta_{t}$. We will maintain this lower envelope using an adapted version of the data structure by Overmars and van Leeuwen [38]. Ideally, we would maintain the lower envelope of $\Delta_{s},\ldots,\Delta_{t}$ directly. However, reassigning a single blue point $b_{j}$ in the matching $\phi$, may cause many functions $\Delta_{k}$ to change. So, we implicitly represent each function $\Delta_{j}$ as a sum of $\Delta^{\prime}_{k}$ functions.
@@
Let $B_{i}$ be a run of size $k$. We can represent the current piece of the lower envelope $g_{i}$ such that we can find the root of (this piece of) $g_{i}$ in $\mathcal{O}(\log k)$ time, and insert or remove any point in $B_{i}$ in $\mathcal{O}(\log^{2}k)$ time.
@@
We will maintain this lower envelope using (a slightly adapted version of) the data structure by Overmars and van Leeuwen [38]. They present a data structure to store a lower envelope of $m$ lines while allowing queries such as line intersections in $\mathcal{O}(\log m)$ time, as well as insertions and deletions in $\mathcal{O}(\log^{2}m)$ time. We could insert the lines $\Delta_{s},\ldots,\Delta_{t}$ into this data structure so we can efficiently find the root of $g_{i}$. However, an update to $\phi$ causes an update to a function $\Delta^{\prime}_{k}$, can cause changes in many functions $\Delta_{j}$, and therefore require significant changes to $g_{i}$. To do handle such updates efficiently, we will make a small modification to the data structure.
As a static lower envelope data structure, they use a binary tree whose leaves are only the lines that contribute to the lower envelope, sorted based on slope. To take advantage of this, we first note that $\Delta_{j}(\tau)$ never has greater slope than $\Delta_{j+1}$, because $\Delta_{j}(\tau)=\Delta^{\prime}_{j}(\tau)+\Delta_{j+1}(\tau)$ and $\Delta^{\prime}_{j}$ has a non-positive slope. This means that in such a tree we can have the leaves sorted based on the indices of the points. To allow for efficient updates, we will not store these functions at the corresponding leaf nodes, but implicitly along the path from the root to that leaf. More concretely, for a node $v$ whose subtree has leftmost leaf $b_{s}$ and where the leaf to the right of its rightmost leaf is $b_{t}$, we store $\ell_{v}(\tau)=\sum_{j=s}^{t-1}\Delta^{\prime}_{j}(\tau)$. If there is no leaf further to the right, then we instead let $b_{t}$ be the rightmost point of the complete run and set $\ell_{v}(\tau)=\sum_{j=s}^{t}\Delta^{\prime}_{j}(\tau)$. Now for a given leaf $b_{j}$, we get exactly $\Delta_{j}$ if we add $\ell_{v}$ for every node $v$ where the path to $b_{j}$ visits the right sibling of $v$.
All queries on the lower envelope work by moving down a path from the root, so we can keep track of the sum accumulated along the path so far and offset the querying line or lower envelope with this. To steer the search path, each node $v$ also stores pointers to its subtree’s leftmost and rightmost leaf. When using these pointers, we need to be careful to still add the appropriate offsets so we actually consider the functions $\Delta_{j}(\tau)$: for the rightmost leaf this offset is zero, for the leftmost leaf $b_{j}$ this is $\ell_{v}(\tau)-\Delta^{\prime}_{j}$. This does not affect query times.
Given the lower envelope for $b_{s},\ldots,b_{j}$ and $b_{j+1},\ldots,b_{t}$, the lower envelope for $b_{s},\ldots,b_{t}$ can be calculated as in the original data structure: by finding where the two envelopes intersect, splitting off the leaves of $b_{s},\ldots,b_{j}$ left from the intersection and of $b_{j+1},\ldots,b_{t}$ right from it, then merging what remains of the two trees. To maintain the correct functions $\ell_{v}(\tau)$, we only need to update the nodes along the path to the rightmost leaf that came from $b_{s},\ldots,b_{j}$. Thus, this operation still takes $\mathcal{O}(\log(t-s))$ time. We can now make the dynamic data structure of Overmars and van Leeuwen [38] use this modified static lower envelope structure to modify it while keeping the same asymptotic running times. ∎
@@
Once we have processed all events, the algorithm has found an optimal translation $\tau^{*}$. We run the sweep once more from the start, and stop at translation $\tau^{*}$, then report the current matching $\phi$ as an optimal matching. Together with Theorem 3.2, this thus establishes Theorem 1.1.

### v2

@@ 3.2 Asymmetric Case
$\mathcal{D}_{B,R}(\phi,\tau)\leq\mathcal{D}_{B,R}(\psi,\tau)$ and $\mathcal{D}_{B,R}(\phi^{\prime},\tau^{\prime})\leq\mathcal{D}_{B,R}(\psi^{\prime},\tau^{\prime}).$ (3.2)
We say that $b\in B$ is a crossing if $\phi^{\prime}(b)<\phi(b)$. Notice that every $b\in C\cap B$ is a crossing. Next, we classify crossings into types based on the order of the points:
  • Type BBRR: $b+\tau<b+\tau^{\prime}\leq\phi^{\prime}(b)<\phi(b)$,
  • Type RRBB: $\phi^{\prime}(b)<\phi(b)\leq b+\tau<b+\tau^{\prime}$,
  • Type RBBR: $\phi^{\prime}(b)\leq b+\tau<b+\tau^{\prime}\leq\phi(b)$,
  • Type BRBR: $b+\tau\leq\phi^{\prime}(b)<b+\tau^{\prime}\leq\phi(b)$,
  • Type RBRB: $\phi^{\prime}(b)\leq b+\tau<\phi(b)\leq b+\tau^{\prime}$,
  • Type BRRB: $b+\tau\leq\phi^{\prime}(b)<\phi(b)\leq b+\tau^{\prime}$.
Figure 3.4: Illustration of crossing types. The left figure shows a BBRR crossing and the right figure shows a RBBR crossing.
Note that for any crossing $b$ of type BBRR or RRBB it holds that:
$|\phi(b)-b-\tau|+|\phi^{\prime}(b)-b-\tau^{\prime}|=|\psi(b)-b-\tau|+|\psi^{\prime}(b)-b-\tau^{\prime}|.$
Moreover, for any crossing $b$ of type RBBR, BRBR, RBRB or BRRB, denoting by $x_{1}\leq x_{2}<x_{3}\leq x_{4}$ the numbers $b+\tau,b+\tau^{\prime},\phi(b)$ and $\phi^{\prime}(b)$ in sorted order, we have:
$|\phi(b)-b-\tau|+|\phi^{\prime}(b)-b-\tau^{\prime}|$ $=x_{3}+x_{4}-x_{1}-x_{2},$
$|\psi(b)-b-\tau|+|\psi^{\prime}(b)-b-\tau^{\prime}|$ $=x_{2}+x_{4}-x_{1}-x_{3}.$
Since $x_{3}>x_{2}$, it follows that for any crossing $b$ of type RBBR, BRBR, RBRB or BRRB we have:
$|\phi(b)-b-\tau|+|\phi^{\prime}(b)-b-\tau^{\prime}|>|\psi(b)-b-\tau|+|\psi^{\prime}(b)-b-\tau^{\prime}|.$
By summing up these inequalities over all $b\in C\cap B$, if at least one crossing in $C$ has type RBBR, BRBR, RBRB or BRRB we obtain
$\mathcal{D}_{B,R}(\phi,\tau)+\mathcal{D}_{B,R}(\phi^{\prime},\tau^{\prime})>\mathcal{D}_{B,R}(\psi,\tau)+\mathcal{D}_{B,R}(\psi^{\prime},\tau^{\prime}),$
contradicting inequalities (3.2).
It remains to consider the case that all crossings in $C$ have type BBRR or RRBB. Note that if the leftmost crossing $b=\min(B\cap C)$ has type BBRR, then we can improve the matching $\phi$ by changing $\phi(b)$ to $\phi^{\prime}(b)$; this contradicts the assumption that $\phi$ is an optimal monotone matching of $\mathrm{EMD}_{p}(B+\tau^{\prime},R)$. Symmetrically, we obtain a contradiction if the rightmost crossing $b=\max(B\cap C)$ has type RRBB. Finally, note that since $C$ is a connected component and both $\phi$ and $\phi^{\prime}$ are monotone, if $C$ has a leftmost crossing of type RRBB and a rightmost crossing of type BBRR, then in between these two there must be at least one crossing of type RBBR, BRBR, RBRB or BRRB. As this case was handled in the previous paragraph, we finished the proof. ∎
We can now come back to breakpoints and runs of $\phi$.
@@
A breakpoint $\tau$ of type (ii) corresponds to a pair of optimal monotonically increasing matchings $\phi,\phi^{\prime}$ for which for all points $b\in B$ we have $\phi(b)\leq\phi^{\prime}(b)$. Furthermore, consider a run $b_{s},\ldots,b_{t}$ of $\phi$ and a point $b_{i}$ with $i\in\{s,\ldots,t\}$. If $\phi(b_{i})<\phi^{\prime}(b_{i})$, then $\phi(b_{j})<\phi^{\prime}(b_{j})$ for all $j\in\{i,\ldots,t\}$.
Proof.
Recall that $\tau$ is a breakpoint of type (ii) if there are two distinct matchings $\phi$ and $\phi^{\prime}$ that both realize the optimal cost at $\tau$. Assume that $\phi$ is also optimal for a translation $\tau^{\prime}<\tau$ or that $\phi^{\prime}$ is also optimal for a translation $\tau^{\prime\prime}>\tau$. Now by Lemma 3.4 we have $\phi(b)\leq\phi^{\prime}(b)$ for every $b\in B$. Next, recall that $b_{s},\dots,b_{t}$ form a run if $\phi$ matches them to consecutive red points. Since $\phi(b_{i+1})$ is matched to the red point right after $\phi(b_{i})$, now $\phi^{\prime}(b_{i})>\phi(b_{i})$ implies in particular that $\phi^{\prime}(b_{i})\geq\phi(b_{i+1})$. With monotonicity this gives us that $\phi^{\prime}(b_{i+1})>\phi^{\prime}(b_{i})\geq\phi(b_{i+1})$. By induction, the same holds for the remaining $j\in\{i+2,\dots,t\}$. ∎
@@
In our sweep line algorithm, we will maintain a current optimal matching $\phi$. At each breakpoint of type (i) we will have an event to update the cost function of the matching. Furthermore, it follows from Corollary 3.8 that when we sweep over a breakpoint of type (ii), we can decompose the changes to the matching using a series of atomic events. In each such atomic event there is some suffix $b_{j},\ldots,b_{t}$ of a run $b_{s},\ldots,b_{t}$ that $\phi$ currently matches to $r_{u-t+j},\dots,r_{u}$ that will become matched to $r_{u-t+j+1},\ldots,r_{u+1}$. As we argued in the proof of Lemma 3.9, the total number of such events is only $\mathcal{O}(nm)$. Next, we express how we can efficiently compute the next such atomic event, and handle it.
Consider a run $B_{i}=b_{s},\ldots,b_{t}$ induced by $\phi$ at time $\tau$. Our aim is to find the smallest $\tau^{\prime}\geq\tau$ at which there is an atomic type (ii) event involving a suffix $b_{j},\ldots,b_{t}$ of $B_{i}$. Hence, for a given suffix $b_{j},\ldots,b_{t}$, we wish to maintain when it starts being beneficial to match $b_{j},\ldots,b_{t}$ to $r_{u-t+j+1},\ldots,r_{u+1}$ rather than to $r_{u-t+j},\ldots,r_{u}$.
@@
Observe that this function is piecewise linear, and non-increasing. Moreover, the breakpoints coincide with type (i) breakpoints of $f$ at which $b+\tau$ coincides with a red point. Hence, in between any two consecutive events, we can consider $\Delta^{\prime}_{j}$ as a linear function. See Figure 3.5 for an illustration.
@@
Let $\tau_{1},\tau^{\prime}]\ni\tau$ be a maximal interval on which $f(\tau)$ is linear, let $\tau^{\prime}$ be a type (ii) breakpoint, and let $\phi$ be an optimal matching for $\tau$. Then there is a run $B_{i}$ induced by $\phi$, and $\tau^{\prime}$ is a root of the function $g_{i}(\tau)$.
@@
At any moment of our sweep, we maintain a single piece of $g_{i}$. Hence, this piece is the lower envelope of a set of linear functions $\Delta_{s},\ldots,\Delta_{t}$. We will maintain this lower envelope using an adapted version of the data structure by Overmars and van Leeuwen [overmars1981]. Ideally, we would maintain the lower envelope of $\Delta_{s},\ldots,\Delta_{t}$ directly. However, reassigning a single blue point $b_{j}$ in the matching $\phi$, may cause many functions $\Delta_{k}$ to change. So, we implicitly represent each function $\Delta_{j}$ as a sum of $\Delta^{\prime}_{k}$ functions.
@@
Let $B_{i}$ be a run of size $k$. Using $\mathcal{O}(k\log k)$ space, we can represent the current piece of the lower envelope $g_{i}$ such that we can find the root of (this piece of) $g_{i}$ in $\mathcal{O}(\log k)$ time, and insert or remove any point in $B_{i}$ in $\mathcal{O}(\log^{2}k)$ time.
@@
We will maintain this lower envelope using (a slightly adapted version of) the data structure by Overmars and van Leeuwen [overmars1981]. They present a data structure to store a lower envelope of $k$ lines while allowing queries such as line intersections in $\mathcal{O}(\log k)$ time, as well as insertions and deletions in $\mathcal{O}(\log^{2}k)$ time. We could insert the lines (representing) $\Delta_{s},\ldots,\Delta_{t}$ into this data structure so we can efficiently find the root of $g_{i}$. However, an update to $\phi$ causes an update to a function $\Delta^{\prime}_{k}$, which in turn could cause changes in many functions $\Delta_{j}$, and could require significant changes to $g_{i}$. We therefore extend the data structure of Overmars and van Leeuwen to support such updates efficiently.
The Overmars and van Leeuwen data structure.
The data structure by Overmars and van Leeuwen is a balanced binary tree, whose leaves store the lines $\Delta_{s},\ldots,\Delta_{t}$ in the decreasing slope order. For each node $v$, let $\Delta^{v}$ denote the (ordered) set of lines stored in the leaves below $v$, and let $E^{v}$ denote the ordered set of lines that appear in left-to-right order on the lower envelope $\mathcal{E}^{v}$ of $\Delta^{v}$. Overmars and van Leeuwen observe that (i) this set of lines $E^{v}$ defining the lower envelope $\mathcal{E}^{v}$ actually forms an ordered subset of the lines in $\Delta^{v}$, i.e. $E^{v}\subseteq\Delta^{v}$, and (ii) that $E^{v}$ consists of a prefix of $E^{\ell}$ concatenated with a suffix of $E^{r}$, where $\ell$ and $r$ are the left and right child of $v$, respectively. See Figure 3.6.
Figure 3.6: The set of lines $E^{v}$ contributing to the lower envelope $\mathcal{E}^{v}$ are ordered by decreasing slope. Furthermore, the lower envelopes $E^{\ell}$ and $E^{r}$ of the left and right child of $v$ intersect in a single point pp. Therefore, $E^{v}$ consists of a prefix of $E^{\ell}$ concatenated with a suffix of $E^{r}$.
The main idea is then to annotate each node $v$ with this ordered set $E^{v}$. Line intersection queries can then be answered in $\mathcal{O}(\log k)$ time by using the $E^{v}$ set of the root. To insert or delete a line $\Delta_{j}$, we follow a root-to-leaf path. While walking back up the search path, we recompute $E^{v}$ from $E^{\ell}$ and $E^{r}$. In particular, by computing the intersection point of $\mathcal{E}^{\ell}$ and $\mathcal{E}^{r}$ and a constant number of split and concatenate operations on the involved ordered sets $E^{v}$. These operations can be implemented in $\mathcal{O}(\log k)$ time, thus leading to an $\mathcal{O}(\log^{2}k)$ update time.
Observe that the total size of all these $E^{v}$ sets is $\mathcal{O}(k\log k)$. Overmars and van Leeuwen reduce the space usage to $\mathcal{O}(k)$ by not explicitly storing the $E^{v}$ sets. Instead, they make sure that every node only stores the lines from $E^{v}$ that were not stored by the parent of $v$, and the relevant sets $E^{v}$ are reconstructed when performing updates. Furthermore, concatenating a prefix of $E^{\ell}$ with a suffix of $E^{r}$ to form $E^{v}$ typically destroys $E^{\ell}$ and $E^{r}$. They carefully describe how and which information to maintain to restore the sets appropriately.
Our data structure.
We make two small, but important, changes to the above data structure.
  • We observe that in our case the (line representing) $\Delta_{j}(\tau)$ never has greater slope than $\Delta_{j+1}$, because $\Delta_{j}(\tau)=\Delta^{\prime}_{j}(\tau)+\Delta_{j+1}(\tau)$ and $\Delta^{\prime}_{j}$ has a non-positive slope. This means that, at any time, the lines $\Delta_{s},\ldots,\Delta_{t}$ are already ordered by decreasing slope. Hence, we can use the indices of the functions to do the routing; i.e. each node $v$ will store the index $s^{v}$ of the rightmost leaf in its left subtree rather than its slope.
  • We extend the data structure so that given a linear function $f$ with non-positive slope and a value $j\in\{s,\ldots,t\}$, we can efficiently add $f$ to all linear functions $\Delta_{j},\ldots,\Delta_{t}$.
    Observe that we are adding $f$ to the lines that already have the smallest slope, so this does not change the overall ordering, and adding $f$ to all functions in an ordered set $\Delta^{v}$ does not change the ordering of the (slopes of the) lines inside $\Delta^{v}$ either. Moreover, the combinatorial structure of its lower envelope $\mathcal{E}^{v}$ remains unchanged: a function $\Delta_{a}$ is the $h^{\mathrm{th}}$ function in $E^{v}$ if and only if $\Delta_{a}+f$ is the $h^{\mathrm{th}}$ function in the lower envelope of $\{\Delta_{c}+f\mid\Delta_{c}\in E^{v}\}$.
    Hence, each node $v$ in our tree will store some additional linear function $f^{v}$ that we still have to add to all functions in the subtree rooted at $v$. As just argued, this allows us to represent the lower envelope $\mathcal{E}^{v}$ corresponding to node $v$ using $E^{v}$ and $f^{v}$.
    We will use the same augmentation in the binary search trees that represent $E^{v}$ themselves; each node $\nu$ will store some linear function $f^{\nu}$ that should still be added to all functions stored in its subtree. We explicitly store these sets $E^{v}$ in persistent red black trees using path copying [driscoll89makin], so that we can still have access to the original sets $E^{\ell}$ and $E^{r}$ after “combining” them into $E^{v}$. This does not affect the running times: we can split, concatenate, and search using these structures in $\mathcal{O}(\log k)$ time, but it increases the total size of our structure to $\mathcal{O}(k\log k)$.
    Using the above representation, we can still compute the intersection of a (query) line with $\mathcal{E}^{v}$ in $\mathcal{O}(\log k)$ time: such a search follows some root-to-leaf path in the tree representing $E^{v}$, so at every step we can locally apply the offset $f^{\nu}$ corresponding to the visited node $\nu$.
    Similarly, inserting or deleting a function still takes $\mathcal{O}(\log^{2}k)$ time. When we visit node $v$, we “push” its offset towards its children. When moving back up the path, we compute the intersection point between two envelopes $\mathcal{E}^{\ell}$ and $\mathcal{E}^{r}$. This involves some simultaneous root-to-leaf path traversal of traversals of the trees representing $E^{\ell}$ and $E^{r}$, so we can again locally apply the offsets involved. Hence, this still takes $\mathcal{O}(\log k)$ time. We then split and concatenate $E^{\ell}$ and $E^{r}$ into $E^{v}$ in additional $\mathcal{O}(\log k)$ time. (Note that this is where we use that every node $\nu$ in the tree representing $E^{v}$ can store its own offset $f^{\nu}$). Hence, the total time required is still $\mathcal{O}(\log^{2}k)$.
    To add $f$ to all functions $\Delta_{j},\ldots,\Delta_{t}$, we simply add $f$ to $f^{r}$, for all $\mathcal{O}(\log k)$ nodes $r$ hanging from the search path to $\Delta_{j}$ (in particular if the search path visits a node $v$ and its left child, we add $f$ to the right child $r$ of $v$). Note that we can determine which child of node $v$ to visit by comparing $j$ against the index $s^{v}$. When we walk back up the path, we recompute $E^{v}$ from $E^{\ell}$ and $E^{r}$ as before. Hence, the total time required is $\mathcal{O}(\log^{2}k)$.
Using the data structure.
We maintain the functions $\Delta_{s},\ldots,\Delta_{t}$ in the above data structure. The lower envelope $\mathcal{E}^{v}$ represented by the root $v$ is then exactly the function $g_{i}$. Hence, we can compute the root of $g_{i}$ by a line-intersection query on $\mathcal{E}^{v}$. To insert a new point $b_{j}$, we insert the linear function $\Delta_{j}$ into the structure, and increment the existing functions $\Delta_{j+1},\ldots,\Delta_{t}$ by $\Delta^{\prime}_{j}$. This takes $\mathcal{O}(\log^{2}k)$ time. Deleting a point is analogous. ∎
@@
Once we have processed all events, the algorithm has found an optimal translation $\tau^{*}$. We run the sweep once more from the start, and stop at translation $\tau^{*}$, then report the current matching $\phi$ as an optimal matching. Together with Theorem 3.3, this thus establishes Theorem $1.$